OpenStudy (anonymous):

multiplying polynomials x(x-2)+(x+1)(x-4) PLEASE HELP

OpenStudy (anonymous):

For the first part, do the distributive property, so multiply x by both the things inside, so x*x=x^2 and x*-2=-2x. For the second part do FOIL (first, ouside, inside, last). So x*x=x^2, -4*x=-4x, 1*x=x, and 1*-4=-4. Then you just add everything together like so: x^2 + (-2x) + x^2 + (-4x) + x + (-4) = x^2 - 2x + x^2 - 4x + x -4 Then combine like terms and you get your final answer: 2x^2 - 5x - 4